OpenStudy (kendricklamar2014):

A true-false test has 5 questions. What is the probability of guessing the correct answers to all the questions? A. 1/25 B. 1/10 C. 1/32 D. 1/7

OpenStudy (kendricklamar2014):

@YanaSidlinskiy

OpenStudy (zubhanwc3):

Disclaimer, i could be wrong. First off, we can easily conclude that its not B or D, so from my memory(i havent done this in forever) its either 5! or 5^2 5! is 180, so thats not it, and so i would choose A can 5^2 is 25

OpenStudy (kendricklamar2014):

Thanks

OpenStudy (zubhanwc3):

please give me a medal if it is correct.

OpenStudy (zarkon):

\[\left(\frac{1}{2}\right)^5\]

OpenStudy (zarkon):

\[=\frac{1}{32}\]

OpenStudy (kendricklamar2014):

Is it A or C?

OpenStudy (zarkon):

each question has a 1 in 2 chance of being correct...so a prob of 1/2 (for each t/f question) if the questions are independent then you multiply the probabilities together... that is \((1/2)(1/2)(1/2)(1/2)(1/2)=(1/2)^5=1/32\)

OpenStudy (kendricklamar2014):

Ok thanks @Zarkon

OpenStudy (zarkon):

This follows a binomial distribution. X being binomial with p=1/2 and n=5 \[P(X=x)={n\choose x}\left(p\right)^{x}\left(1-p\right)^{n-x}\] then \[P(X=5)={5\choose 5}\left(\frac{1}{2}\right)^{5}\left(1-\frac{1}{2}\right)^{5-5}=\frac{1}{32}\]

OpenStudy (kendricklamar2014):

thanks for everything :)